# Assurance of AI Systems From a Dependability Perspective

### v1

@@ Abstract
We apply our perspective to autonomous systems, AI systems for specific functions, generic AI such as Large Language Models, and to Artificial General Intelligence (AGI), and we propose current best practice and an agenda for research.
@@ 1 Introduction: Assurance for Traditional Systems
“If a builder build a house for some one, and does not construct it properly, and the house which he built fall in and kill its owner, then that builder shall be put to death” [67, Section 229].
@@
The dependability perspective may be contrasted with what we call the trustworthiness perspective, which does claim some assurance in the behavior of AI and ML components that have been developed, analyzed, tested, augmented, or restricted in various ways. Both perspectives have merit and in practice there is a continuum or spectrum between them. In particular, we envisage guarded architectures that are recursively structured where “first level” guards might use some AI and ML and themselves be guarded by simpler systems, eventually bottoming out on conventionally engineered and assured guards so that the overall architecture provides “defense in depth.” The architecture and its assurance will vary according to how much assurance “credit” is taken for trustworthiness or diversity of AI and ML components [25]. We call this the dependability/trustworthiness spectrum; a “pure” dependability perspective takes no credit for trustworthiness of AI and ML components, while a “pure” trustworthiness perspective claims full assurance credit for those components.
In the remainder of this section, we describe the dependability perspective on assurance for traditional systems that do not employ AI or ML. Subsequent sections introduce increasing amounts of AI and ML and we discuss approaches and concerns regarding their assurance from perspectives toward the dependability end of the spectrum. We invite others to provide complementary studies toward its trustworthiness end.
@@ 1.1 Traditional Systems and their Assurance
Requirements concern what the system will do, not how it will do it, so they should largely be described in terms of changes the system is to bring about in the environment (this is a key insight due to Michael Jackson [80]). How the system will do its task is developed in specifications for the defined behavior of the system and the architecture of its components. Architecture is a generalization of partitioning (often portrayed by “boxes and arrows” diagrams) and its purposes are to identify fault containment regions that limit fault propagation among components, to identify critical components and limit their complexity (because complexity is a source of faults), and generally organize things so that dependability relies on only the architecture and the defined behavior of the critical components [29].
@@ 1.2 From Assurance to Dependability
Cars, most of whose accidents are caused by human driver error, are among the most dangerous consumer goods, with about 35,000 deaths per year in the United States and a fatal accident rate of a little over 10 per billion miles. It is intended that self-driving cars should be safer than human drivers, so it might seem reasonable to set the target at no more than 1 death per billion miles, which is $4\times 10^{-7}$ per hour, if we assume 25 mph. However, that is a rather technical assessment that pays no attention to likely public reaction. Herbert Diess, the CEO of Volkswagen is quoted on their website with a more realistic assessment: “A ratio of ten-to-one is nowhere near good enough. We have approximately 3,200 traffic fatalities in Germany each year. It would be a disaster if we had even 320 deaths due to driverless cars.” [^5] In the USA, a ten-to-one ratio would yield 3,500 fatalities due to driverless cars, which would surely also be unacceptable. Thus, it is widely accepted that the safety target for self-driving cars should be 100 or even 1,000 times better than human drivers, [^6] This sensitivity is illustrated by a recent accident in San Francisco: a car with a human driver hit a pedestrian and threw them under the wheels of a Cruise self-driving car, which then dragged the victim for several yards [131]. Subsequently, Cruise lost their license to operate in San Francisco (although their poor initial response to the incident probably contributed to this harsh reaction). which brings us into, or even beyond, the requirements for commercial aircraft. But commercial aircraft do get assured and certified and their safety record justifies this, so why is $10^{-9}$ seen as such a challenge?
The answer is that we do not assure aircraft solely by testing. If we test a system for $n$ hours and observe no failures, then in the absence of other information, the best prediction we can make is that the likelihood of no failures in another $n$ hours is about 50–50 [109, p. 73, and sidebar on p. 74]. Hence, to secure assurance for failure rates of $10^{-9}$ per hour we would need to test the system for about $10^{9}$ hours, or around 115,000 years. Even with 1,000 copies of the system on test, this is still well over 100 years of continuous operation and is completely infeasible [35, 109, 87].
@@
We can now see that traditional approaches to assurance, such as described for OP, give us justifiably strong confidence that the system satisfies its critical properties; this can then be augmented by testing and operational experience to deliver (via CBI and Bootstrapping) justified confidence that the system satisfies its dependability goals. In the following sections, we consider systems with AI and ML components that do not lend themselves to traditional methods of assurance and may therefore be ineligible for CBI, and we explore how, and to what extent, dependability can nonetheless be achieved and guaranteed.
@@ 2.1 Runtime Verification
To investigate runtime verification for systems with AI and ML components, we need some general understanding of the likely overall system architecture and the properties that will be checked. The top-level structure of almost any system that employs AI or ML follows from a single insight, which is that any entity that interacts effectively with some aspect of the world must have a model of that aspect of the world [43]. In particular, cyber-physical systems are always based on a model of the controlled “plant” and its environment [57]. For pure control systems, the model may be a collection of differential equations; for systems that involve procedures it may add state machines; and for full CPS it will include integrated formalisms such as timed and hybrid automata. [^9] A model can be thought of as a simulation (or “digital twin”) of the plant and its environment; “model based system/software engineering” (MBSE) builds and experiments with explicit simulations of these plus the intended system using tools such as StateFlow/Simulink and derives the implementation (sometimes automatically) from the system simulation. A criticism of this approach is that the focus on simulation introduces implementation concerns too early into the process (it is difficult to simulate requirements stated as constraints unless theorem proving is used, so specifications are often substituted for requirements) and thereby compromises dependability requirements validation.
@@
An example of the first class is an unmanned aircraft controlled by AI and ML that is deemed safe as long as it remains within some specified portion of airspace (a “geofence”). This is a constraint that can be checked by a conventional navigation system and enforced using a conventional guidance system as an override when violations are detected [46]. Many robots can be guarded by “virtual cages” of this kind. Notice also that when formulating the initial “concept of operation” for a system, it may be possible to adjust the concept so that assurable guards become feasible. For example, rather than an autonomous shuttle bus sharing its route with other vehicles and pedestrians, it could use a dedicated track, or rails, and would not then require a sophisticated AI-based perception system.
An example of the second class is a self-driving car where the action guard checks AI-generated actions for safety and other dependability properties against a model of its environment—i.e., location of other road users and pedestrians, interpretation of traffic signs and markings, detection of road layout, etc. To verify the guard we need assurance for accuracy of the model, and to do this in the traditional way we would want to see its requirements (i.e., what it should represent and what it means for it to do so with acceptable accuracy), its hazards (e.g., inaccuracies that can lead to harm), together with specifications and an implementation that can be shown to achieve the requirements and mitigate the hazards. Typically, none of this is present for models constructed by an AI perception system that uses ML to interpret its sensors, although recent developments are starting to explore some of these topics [74, 37, 24].
@@ 2.2 The Challenge of Assuring Perception
A possible compromise is for the action guard to use a simpler, conservative model constructed by conventional and highly assured software. This is sometimes seen in self-driving cars where the guard is a simple system for automated collision detection and emergency braking, similar to the Automated Driver Assistance System (ADAS) that is provided for some human-driven cars. [^11] The US National Highway Traffic Safety Administration recently finalized a rule (FMVSS No. 127) that will require Automatic Emergency Braking (AEB), to be standard on all passenger cars and light trucks by September 2029. Tests by the American Automobile Association found that current AEBs (which are built to a lower, voluntary standard) are somewhat effective against rear-end collisions, but not at all against sideways collisions at intersections. However, this does not provide full assurance because forward collisions are not the only hazards (Mehmed and colleagues cite data from a NHTSA study of 5.9 million human-driver accidents classified in 37 categories [118]), nor are human accidents necessarily good models for failures of an autonomous system, and nor is emergency braking attractive as the sole method of hazard mitigation. Furthermore, accidents and collisions are not the only hazards that should be considered: for example, the City of San Francisco has reported dozens of incidents where robotaxis interfered with emergency responders, and there must be many other circumstances where self-driving cars increase risk or inconvenience for others without themselves being involved in a collision.
@@
The hazard most generally recognized in perception using ML is lack of “robustness,” meaning that small changes in sensor data may cause its ML-generated interpretation to change abruptly. This concern is validated by so-called “adversarial examples” [161]. Typically demonstrated on image classifiers, the examples are deliberately constructed minor modifications to an input image that are indiscernible to a human observer but cause an ML classifier to change its output, often drastically and inappropriately. Image masks have been developed that will cause misclassification when overlayed on any image input to a given classifier, and there are universal examples that will disrupt any classifier [120]. Furthermore, there are patterns that can be applied to real-world artifacts (e.g., small images that can be stuck to traffic signs) that will cause them to be misread by an object classifier [33].
@@
Accurate perception for world models in general depends not only on robust interpretation of individual sensors but also on effective sensor fusion. This is illustrated by the fatal accident between an Uber self-driving car and a pedestrian walking a bicycle in Arizona on 18th March 2018 [121]. The Uber car used three sensor systems (cameras, radars, and lidar) and fused them using a priority scheme that delivered a “flickering” identification of the victim as the sensor systems’ own classifiers changed their identifications, and as fusion preferred first one sensor system, then another, as listed below [121, Table 1].
@@ 2.3 Assurance through Diversity and Defense in Depth
The topic of assurance through diversity is large and somewhat contentious. There is little doubt that architectures employing diverse components are generally more reliable than single threads. In particular, there is evidence that “portfolio” or “ensemble” perception systems are more reliable than their individual constituents [84]. The difficulty is in demonstrating that diversity provides benefit in any particular case, and in estimating how much benefit it provides [107]. In particular, there is no feasible way to validate failure independence (it is a variant on the infeasibility of assurance by testing), nor strong reasons for believing it. This is because some circumstances are just plain hard to interpret and it is possible that all components may then fail together: consider the scenario with the Cruise self-driving car described earlier—would any training set have included pedestrians being thrown under the wheels? Furthermore, these difficult cases do not “thin out” as more are considered: the distributions seem to have “fat tails” [94]. Thus, the “multiplicative” argument for assurance by diversity is indeed naive.
However, although diversity alone cannot provide strong assurance, it can provide a useful step in a “ladder” of assurance. Nuclear power generation usually employs such a ladder of protection systems providing defense in depth: there is the operational control system, designed to manage the plant efficiently and safely, then a (safety) limitation system that can intervene to ensure the plant behavior stays operational but within some safe envelope, and finally a shutdown system that functions as an assured guard that guarantees to initiate a safe shutdown. [^13] A complete shutdown can be a complex operation involving both automated initiation and intervention by expert operators, depending on the state of the plant and reasons underlying the need to shutdown (cf. the Fukushima incident). The operational and limitation systems are carefully engineered but do not guarantee the dependability goals established for nuclear power: that is accomplished by the assured shutdown system. But the operational and limitation systems are diverse in function and construction and although this does not provide strong assurance, the presence of the limitation system almost certainly reduces demands on the shutdown system. This is beneficial because an emergency shutdown is disruptive and expensive.
@@
One possible mitigation for this hazard is to provide the safety-focused system with separate sensors and perception; alternatively, it could use the same sensors as the primary system but with diverse perception software. Either arrangement would provide the primary and safety-focused systems with different models of the world (e.g., [68]) but it is debatable whether diverse models are manageable without false alarms and whether the ML-generated model for the safety system could be any more assurable than that for the primary. [^14] The model for the safety system could be simpler than that for the primary system. For example, the primary system in a self-driving car must detect objects and their type (car, bicycle, pedestrian etc.) because the action system needs to project their likely motion and future position, whereas the safety system could just use objects and surround them with a conservative “safety box.” Intuitively, we might think that the simpler model of the safety system could be more trustworthy, but it would be difficult to provide credible assurance for this using current methods.
@@
A dual-process architecture with System 1 using PP to integrate diverse perception systems and with System 2 providing diverse perception refinement and action guards is an attractive arrangement for autonomous systems. Although assurance for its AI and ML components will be lacking, its overall architecture has a rational structure that can justify modest assurance based on redundancy and diversity. Modest assurance is insufficient for system dependability, but that is not the claim it needs to support: instead, contributing to defense in depth, it supports the claim that it reduces demands on the traditionally engineered emergency backup to a tolerable level, [^16] For example, the primary system of a self-driving car might see an object ahead and classify it as a cardboard box and be prepared to drive over it; as it comes closer, the emergency backup will see it as an unclassified obstruction and slam the brakes on; but the safety system will also have seen it as an unclassified obstruction and could have changed lanes to avoid it. and dependability is assured by the backup.

### v2

@@ Abstract
We apply our perspective to autonomous systems, AI systems for specific functions, general-purpose AI such as Large Language Models, and to Artificial General Intelligence (AGI), and we propose current best practice and an agenda for research.
@@ 1 Introduction: Assurance for Traditional Systems
“If a builder build a house for some one, and does not construct it properly, and the house which he built fall in and kill its owner, then that builder shall be put to death” [69, Section 229].
@@
The dependability perspective may be contrasted with what we call the trustworthiness perspective, which does claim some assurance in the behavior of AI and ML components that have been developed, analyzed, tested, augmented, or restricted in various ways. [^1] Terminology across different fields is always difficult and sometimes contentious; the field of dependability regards “trustworthiness” as a synonym for “dependability,” whereas AI tends to use “safety” as the required property (e.g., the UK Government has an AI Safety Institute) and “trustworthiness” as the means for achieving and assuring it. The adjective “trustworthy” carries a somewhat anthropomorphic tone that we discuss in Section 4 and so we generally prefer the more neutral, engineering terminology of dependability. However, we need a name for the alternative perspective and we use “trustworthy” for that purpose. Both perspectives have merit and in practice there is a continuum or spectrum between them. In particular, we envisage guarded architectures that are recursively structured where “first level” guards might use some AI and ML and themselves be guarded by simpler systems, eventually bottoming out on conventionally engineered and assured guards so that the overall architecture provides “defense in depth.” The architecture and its assurance will vary according to how much assurance “credit” is taken for trustworthiness or diversity of AI and ML components [26]. We call this the dependability/trustworthiness spectrum; a “pure” dependability perspective takes no credit for trustworthiness of AI and ML components, while a “pure” trustworthiness perspective claims full assurance credit for those components.
In the remainder of this section, we describe the dependability perspective on assurance for traditional systems that do not employ AI or ML. Subsequent sections introduce increasing amounts of AI and ML and we discuss approaches and concerns regarding their assurance from perspectives toward the dependability end of the spectrum. We invite others to provide complementary studies toward its trustworthiness end. We stress that the purpose of assurance as we present it is not to impose a brake or burden on development, but to support innovation by anticipating downstream hazards and suggesting creative ways to mitigate them.
@@ 1.1 Traditional Systems and their Assurance
Requirements concern what the system will do, not how it will do it, so they should largely be described in terms of changes the system is to bring about in the environment (this is a key insight due to Michael Jackson [82]). How the system will do its task is developed in specifications for the defined behavior of the system and the architecture of its components. Architecture is a generalization of partitioning (often portrayed by “boxes and arrows” diagrams) and its purposes are to identify fault containment regions that limit fault propagation among components, to identify critical components and limit their complexity (because complexity is a source of faults and also makes it more difficult to discover what faults may be present), and generally organize things so that dependability relies on only the architecture and the defined behavior of the critical components [30].
@@ 1.2 From Assurance to Dependability
Cars are among the most dangerous consumer goods with about 40,000 deaths per year in the United States and a fatal accident rate of a little over 10 per billion miles. It is intended that self-driving cars should be safer than human drivers, so it might seem reasonable (even though driving error is not the only cause of accidents) to set the target at no more than 1 death per billion miles, which is around $10^{-7}$ per hour, given an average speed of 30 mph. However, that is a rather technical assessment that pays no attention to likely public reaction. Herbert Diess, the former CEO of Volkswagen is quoted on their website with a more realistic assessment: “A ratio of ten-to-one is nowhere near good enough. We have approximately 3,200 traffic fatalities in Germany each year. It would be a disaster if we had even 320 deaths due to driverless cars.” Thus, it is widely accepted that the safety target for self-driving cars should be 100 or even 1,000 times better than human drivers, [^6] This sensitivity is illustrated by a recent accident in San Francisco: a car with a human driver hit a pedestrian and threw them under the wheels of a Cruise self-driving taxi, which then dragged the victim for several yards. Subsequently, Cruise lost their license to operate in San Francisco (although their poor initial response to the incident probably contributed to this harsh reaction) [97]. which brings us into, or even beyond, the requirements for commercial aircraft. But commercial aircraft do get assured and certified and their safety record justifies this, so why is $10^{-9}$ seen as such a challenge?
The answer is that we do not assure aircraft solely by testing. If we test a system for $n$ hours and observe no failures, then in the absence of other information, the best prediction we can make is that the likelihood of no failures in another $n$ hours is about 50–50 [113, p. 73, and sidebar on p. 74]. Hence, to secure assurance for failure rates of $10^{-9}$ per hour we would need to test the system for about $10^{9}$ hours, or around 115,000 years. Even with 1,000 copies of the system on test, this is still well over 100 years of continuous operation and is completely infeasible [37, 113, 90].
@@
But suppose experience in operation does reveal a failure, hopefully not catastrophic (in commercial aircraft, for example, it is required that no catastrophic failure may be caused by a single fault). Any such failure reveals an unanticipated fault, arrival rate, circumstance, or hazard, and these could be precursors to a catastrophic failure. Consequently, commercial airplanes operate in a legal and ethical framework where all incidents and accidents are promptly reported and dispassionately investigated. The FAA issues Airworthiness Directives mandating workarounds or corrections to detected faults; in extreme cases it may temporarily ground the fleet (as it did for the 737 MAX in 2019–2020). Bishop [21] constructs a statistical model for this scenario and shows that, under plausible assumptions, detection and repair of faults significantly increases long run safety, even if the fleet continues to operate after a fault has been discovered, and even if repairs may be imperfect. It follows that there is much value in monitoring, analyzing and, if warranted, correcting all non-trivial failures and their precursors.
We can now see that traditional approaches to assurance, such as described for OP, give us justifiably strong confidence that the system satisfies its critical properties; this can then be augmented by testing and operational experience to deliver (via CBI and Bootstrapping) justified confidence that the system satisfies its dependability goals, and this can be reinforced by monitoring during operation. In the following sections, we consider systems with AI and ML components that do not lend themselves to traditional methods of assurance and may therefore be ineligible for CBI, and we explore how, and to what extent, dependability can nonetheless be achieved and guaranteed.
@@ 2.1 Runtime Verification
To investigate runtime verification for systems with AI and ML components, we need some general understanding of the likely overall system architecture and the properties that will be checked. The top-level structure of almost any system that employs AI or ML follows from a single insight, which is that any entity that interacts effectively with some aspect of the world must have a model of that aspect of the world [45]. In particular, cyber-physical systems are always based on a model of the controlled “plant” and its environment [59]. For pure control systems, the model may be a collection of differential equations; for systems that involve procedures it may add state machines; and for full CPS it will include integrated formalisms such as timed and hybrid automata. [^9] A model can be thought of as a simulation (or “digital twin”) of the plant and its environment; “model based system/software engineering” (MBSE) builds and experiments with explicit simulations of these plus the intended system using tools such as StateFlow/Simulink and derives the requirements and specification, and sometimes the implementation (and sometimes does that automatically), from the system simulation. A criticism of this approach is that the focus on simulation introduces implementation concerns too early into the process (it is difficult to simulate requirements stated as constraints unless theorem proving is used, so specifications are often substituted for requirements) and thereby compromises dependability requirements validation.
@@
An example of the first class is an unmanned aircraft controlled by AI and ML that is deemed safe as long as it remains within some specified portion of airspace (a “geofence”). This is a constraint that can be checked by a conventional navigation system and enforced using a conventional guidance system as an override when violations are detected [48]. Many robots can be guarded by “virtual cages” of this kind. Notice also that when formulating the initial “concept of operation” for a system, it may be possible to adjust the concept so that assurable guards become feasible. For example, rather than an autonomous shuttle bus sharing its route with other vehicles and pedestrians, it could use a dedicated and isolated track, and would not then require a sophisticated AI-based perception system.
An example of the second class is a self-driving car where the action guard checks AI-generated actions for safety and other dependability properties against a model of its environment—i.e., location of other road users and pedestrians, interpretation of traffic signs and markings, detection of road layout, etc. To verify the guard we need assurance for accuracy of its model, and to do this in the traditional way we would want to see its requirements (i.e., what it should represent and what it means for it to do so with acceptable accuracy), its hazards (e.g., inaccuracies that can lead to harm), together with specifications and an implementation that can be shown to achieve the requirements and mitigate the hazards. Typically, none of this is present for models constructed by an AI perception system that uses ML to interpret its sensors, although recent developments are starting to explore some of these topics [76, 39, 25].
@@ 2.2 The Challenge of Assuring Perception
A possible compromise is for the action guard to use a simpler, conservative model constructed by conventional and highly assured software. This is sometimes seen in self-driving cars where the guard is a simple system for automated collision detection and emergency braking, similar to the Automated Driver Assistance System (ADAS) that is provided for some human-driven cars. [^11] The US National Highway Traffic Safety Administration recently finalized a rule (FMVSS No. 127) that will require Automatic Emergency Braking (AEB), to be standard on all passenger cars and light trucks by September 2029. Tests by the American Automobile Association found that current AEBs (which are built to a lower, voluntary standard) are somewhat effective against rear-end collisions, but not at all effective against sideways collisions at intersections. However, this does not provide full assurance because forward collisions are not the only hazards (Mehmed and colleagues cite data from a NHTSA study of 5.9 million human-driver accidents classified in 37 categories [122]), nor are human accidents necessarily good models for failures of an autonomous system, and nor is emergency braking attractive as the sole method of hazard mitigation. And notice that to prevent excessive activation of the emergency system, the primary system of a self-driving car must take its behavior and capabilities into account. For example, the primary action subsystem may calculate (correctly) that a certain maneuver is optimal and safe, but that it is also likely to activate the more hair-trigger response of the emergency system, and so it must choose a different behavior. In general, systems employing defense in depth must be designed and developed as a whole to ensure that unintended differences among the layers do not cause unnecessary loss of availability, while safety arguments based on diversity are nevertheless preserved. These are challenging requirements.
Furthermore, accidents and collisions are not the only hazards that should be considered: for example, the City of San Francisco has reported dozens of incidents where robotaxis interfered with emergency responders, and there must be many other circumstances where self-driving cars increase risk or inconvenience for others without themselves being involved in a collision.
@@
The hazard most generally recognized in perception using ML is lack of “robustness,” meaning that small changes in sensor data may cause its ML-generated interpretation to change abruptly. This concern is validated by so-called “adversarial examples” [165]. Typically demonstrated on image classifiers, the examples are deliberately constructed minor modifications to an input image that are indiscernible to a human observer but cause an ML classifier to change its output, often drastically and inappropriately. Image masks have been developed that will cause misclassification when overlayed on any image input to a given classifier, and there are universal examples that will disrupt any classifier [124]. Furthermore, there are patterns that can be applied to real-world artifacts (e.g., small images that can be stuck to traffic signs) that will cause them to be misread by an image classifier [35].
@@
Accurate perception for world models in general depends not only on robust interpretation of individual sensors but also on effective sensor fusion. This is illustrated by the fatal accident between an Uber self-driving car and a pedestrian walking a bicycle in Arizona on 18th March 2018 [125]. The Uber car used three sensor systems (cameras, radars, and lidar) and fused them using a priority scheme that delivered a “flickering” identification of the victim as the sensor systems’ own classifiers changed their identifications, and as fusion preferred first one sensor system, then another, as listed below [125, Table 1].
@@ 2.3 Assurance through Diversity and Defense in Depth
The topic of assurance through diversity is large and somewhat contentious. There is little doubt that architectures employing diverse components are generally more reliable than single threads. In particular, there is evidence that “portfolio” or “ensemble” perception systems are more reliable than their individual constituents [87]. The difficulty is in demonstrating that diversity provides benefit in any particular case, and in estimating how much benefit it provides [111]. In particular, there is no feasible way to validate failure independence (it is a variant on the infeasibility of assurance by testing), [^13] However, it is feasible to validate modest degrees of independence, and indications of its absence. nor strong reasons for believing it. This is because some circumstances are just plain hard to interpret and it is possible that all components may then fail together: consider the scenario with the Cruise self-driving taxi described in Footnote 6—would any training set have included pedestrians being thrown under the wheels? Furthermore, these difficult cases do not “thin out” as more are considered: the distributions seem to have “fat tails” [98]. Thus, the “multiplicative” argument for assurance by diversity is indeed naive.
However, although diversity alone cannot provide strong assurance, it can provide a useful step in a “ladder” of assurance. Nuclear power generation usually employs such a ladder of protection systems providing defense in depth: there is the operational control system, designed to manage the plant efficiently and safely, then a (safety) limitation system that can intervene to ensure the plant behavior stays operational but within some safe envelope, and finally a shutdown system that functions as an assured guard that guarantees to initiate a safe shutdown when safety parameters are violated. [^14] A complete shutdown can be a complex operation involving both automated initiation and intervention by expert operators, depending on the state of the plant and reasons underlying the need to shutdown (cf. the Fukushima incident). The operational and limitation systems are carefully engineered but do not guarantee the dependability goals established for nuclear power: that is accomplished by the assured shutdown system. But the operational and limitation systems are diverse in function and construction and although this does not provide strong assurance, the presence of the limitation system almost certainly reduces demands on the shutdown system. This is beneficial because an emergency shutdown is disruptive and expensive.
@@
One possible mitigation for this hazard is to provide the safety-focused system with separate sensors and perception; alternatively, it could use the same sensors as the primary system but with diverse perception software. Either arrangement would provide the primary and safety-focused systems with different models of the world (e.g., [70]) but it is debatable whether diverse models are manageable without false alarms and whether the ML-generated model for the safety system could be any more assurable than that for the primary. [^15] The model for the safety system could be simpler than that for the primary system. For example, the primary system in a self-driving car must not only detect objects, but their type (car, bicycle, pedestrian etc.) and fine details such as where drivers and pedestrians are looking, because the action system needs to project their likely motion and future position, whereas the safety system could just use objects and surround them with a conservative “safety box.” Intuitively, we might think that the simpler model of the safety system could be more trustworthy, but it would be difficult to provide credible assurance for this using current methods.
@@
A dual-process architecture with System 1 using PP to integrate diverse perception systems and with System 2 providing diverse perception refinement and action guards is an attractive arrangement for autonomous systems. Although assurance for its AI and ML components will be lacking, its overall architecture has a rational structure that can justify modest assurance based on redundancy and diversity. Modest assurance is insufficient for system dependability, but that is not the claim it needs to support: instead, contributing to defense in depth, it supports the claim that it reduces demands on the traditionally engineered emergency backup to a tolerable level, [^17] For example, the primary system of a self-driving car might see an object ahead and classify it as a cardboard box and be prepared to drive over it; as it comes closer, the emergency backup will see it as an unclassified obstruction and slam the brakes on; but the safety system will also have seen it as an unclassified obstruction and could have changed lanes to avoid it “just in case.” On a larger scale, the safety system might veto a time-optimal route selection by the primary system because it considers it unduly hazardous (e.g., icy). and dependability is assured by the backup.